# Preference for Verifiability

### v1

@@ Appendix C Proof of Theorem 2
$\alpha a\oplus(1-\alpha)c$ $\succsim\alpha b\oplus(1-\alpha)c\quad$ $\Leftrightarrow^{def.}$ $\quad\alpha a\oplus(1-\alpha)c$ $\succsim^{\dagger}\alpha b\oplus(1-\alpha)c$ (C.9)
The next lemma is trivial.
Lemma 5 (Supermodularity-Submodularity Correspondence).
[…]
Lemma 6 (Critical Event Modularity Correspondence).
@@
$U(a)=$ $-\int\max_{E\in\mathscr{V}}\min_{s\in E}u(a(s))d\mu$
$=$ $-\int\max_{E\in\mathscr{V}}-\max_{s\in E}-u(a(s))d\mu$
$=$ $\int\min_{E\in\mathscr{V}}\max_{s\in E}-u(a(s))d\mu$ (C.10)
Appendix D Proof of Proposition 2
This result is a standard result. The only interesting question is whether the definition of comparative risk preference is compatible with different verifiable events for each decision maker. Note that the definition of comparative risk preference only depends on preferences over constant acts. Further, both in expected verification utility and expected obfuscation utility the definition of the constant act $1/2x\oplus 1/2y$ does not depend on the verifiable events. Thus, risk attitudes are comparable across decision makers with different verifiable events and across decision makers with obfuscation and verification utility.
Appendix E Proof of Proposition 3
@@
$1\Rightarrow 2:$ If $\gamma E\beta\sim^{2}\gamma(E-F)\beta,$ then $E$ is not critical. Since the critical events of decision maker 2 contain all critical events of decision maker 1, $E$ cannot be critical for decision maker 1. Thus, $\gamma E\beta\sim^{1}\gamma(E-F)\beta$.
$2\Rightarrow 1:$ If $E$ is critical for decision maker 1 but not critical for decision maker 2, then for decision maker $2$ there exists some set $F\subset E$ such that $\gamma E\beta\sim^{2}\gamma(E-F)\beta$. But then also $\gamma E\beta\sim^{1}\gamma(E-F)\beta$ and thus $E$ cannot be critical for decision maker 1. Thus, the critical sets of decision maker 1 are a subset of the critical sets of decision maker 2. ∎
Appendix F Proofs of Corollaries 1-3
@@
For the proof of Corollary 1, consider a decision problem $A$ in which the loss given a verification-seeking decision maker is nonzero. Then there exists an event $E$ such that no proper subset is verifiable. In particular, let $F$ and $G$ be disjoint proper subsets of $E$. Consider the acts $\gamma F\beta G\delta$ and $\gamma G\beta F\delta^{\prime}$ with $\gamma\succ\beta$. Assume that the utility difference between $\delta$ and $\delta^{\prime}$ is small. An expected utility maximizing DM strictly prefers the former to the latter if and only if $\mu(F)>\mu(G)$. A verification seeking decision maker strictly prefers the former to the latter if and only if $\beta\succ\delta\succ\delta^{\prime}$. An obfuscation seeking decision maker strictly prefers the former to the latter if and only if $\delta\succ\delta^{\prime}\succ\gamma$. Thus, it is possible that the expected utility maximizer agrees on the optimal action with the verification seeker or the obfuscation seeker without agreeing with the other.
The proof of 2 is trivial, the expected utility of the maximizer of $U$ can at most be the maximum expected utility on a decision problem.
@@
Appendix G Axioms for Biseparable Preferences
In the main text, a biseparable preference is assumed. Here we restate the axioms used by [29] to characterize biseparable preferences. The axioms are slightly adjusted to fit the notation of the present paper.
@@ Appendix G Axioms for Biseparable Preferences
Continuity ensures the existence of certainty equivalents and therefore the subjective mixtures [38, 30, 31] are well defined:
@@
$xEy\succsim x^{\prime}Ey^{\prime}\quad\Rightarrow\quad(xEy)\oplus_{B}(x^{\prime\prime}Ey^{\prime\prime})\succsim(x^{\prime}Ey^{\prime})\oplus_{B}(x^{\prime\prime}Ey^{\prime\prime}).$ (G.11)
Theorem 3 (Biseparable Utility Representation [29]).

### v2

@@ Appendix E Proof of Theorem 2
$\alpha a\oplus(1-\alpha)c$ $\succsim\alpha b\oplus(1-\alpha)c\quad$ $\Leftrightarrow^{def.}$ $\quad\alpha a\oplus(1-\alpha)c$ $\succsim^{\dagger}\alpha b\oplus(1-\alpha)c$ (E.23)
Lemma 8 (Proof Axiom Correspondence).
@@
[…]
$U(a)=$ $-\int_{s\in\mathscr{S}^{*}}\max_{E\in\mathscr{V}:s\in E}\min_{s\in E}u(a(s))d\mu$
$=$ $-\int_{s\in\mathscr{S}^{*}}\max_{E\in\mathscr{V}:s\in E}-\max_{s\in E}-u(a(s))d\mu$
$=$ $\int_{s\in\mathscr{S}^{*}}\min_{E\in\mathscr{V}:s\in E}\max_{s\in E}-u(a(s))d\mu$ (E.24)
Redefining the utility over consequences yields the desired result. ∎
Appendix F Proof of Proposition 2
This result is a standard result. The only interesting question is whether the definition of comparative risk preference is compatible with different verifiable events for each decision maker. Note that the definition of comparative risk preference only depends on preferences over constant acts. Further, both in expected verification utility and expected obfuscation utility the definition of the constant act $1/2x\oplus 1/2y$ does not depend on the verifiable events. Thus, risk attitudes are comparable across decision makers with different verifiable events and across decision makers with obfuscation and certification utility.
Appendix G Proof of Proposition 3
@@
We prove the result for expected certification utility, the derivation for expected obfuscation utility is analogous.
$1\Rightarrow 2:$ If for a nonnull $F$, $\gamma E\beta\sim^{2}\gamma(E-F)\beta,$ then $E$ is not sensitive for decision maker 2. Since the sensitive events of decision maker 2 contain all sensitive events of decision maker 1, $E$ cannot be sensitive for decision maker 1. Thus, $\gamma E\beta\sim^{1}\gamma(E-F)\beta$.
$2\Rightarrow 1:$ If $E$ is sensitive for decision maker 1 but not sensitive for decision maker 2, then for decision maker $2$ there exists some nonnull event $F\subset E$ such that $\gamma E\beta\sim^{2}\gamma(E-F)\beta$. But then also $\gamma E\beta\sim^{1}\gamma(E-F)\beta$ and thus $E$ cannot be sensitive for decision maker 1. Thus, the sensitive sets of decision maker 1 are a subset of the sensitive sets of decision maker 2.
Appendix H Proofs of Corollaries 3-5
@@
For the proof of Corollary 3, consider a decision problem $A$ in which the loss given a certification-seeking decision maker is nonzero. Then there exists an event $E$ such that no subset is verifiable but a superset of which is verifiable. In particular, let $F$ and $G$ be disjoint proper subsets of $E$. Consider the acts $\gamma F\beta G\delta$ and $\gamma G\beta F\delta^{\prime}$ with $\gamma\succ\beta$. Assume that the utility difference between $\delta$ and $\delta^{\prime}$ is small. An expected utility maximizing DM strictly prefers the former to the latter if and only if $\mu(F)>\mu(G)$ (up to the small utility difference between $\delta$ and $\delta^{\prime}$). A certification seeking decision maker strictly prefers the former to the latter if and only if $\beta\succ\delta\succ\delta^{\prime}$. An obfuscation seeking decision maker strictly prefers the former to the latter if and only if $\delta\succ\delta^{\prime}\succ\gamma$. Thus, it is possible that the expected utility maximizer agrees on the optimal action with the verification seeker or the obfuscation seeker without agreeing with the other.
The proof of 4 is trivial, the expected utility of the maximizer of $U$ can at most be the maximum expected utility on a decision problem. Since if every event is verifiable the decision maker maximizes expected utility, the result follows.
@@
Appendix I Axioms for Biseparable Preferences
In the main text, a biseparable preference is assumed. Here we restate the axioms used by [undefad] to characterize biseparable preferences. The axioms are slightly adjusted to fit the notation of the present paper.
@@ Appendix I Axioms for Biseparable Preferences
Continuity ensures the existence of certainty equivalents and therefore the subjective mixtures [undefao, undefae, undefaf] are well defined:
@@
$xEy\succsim x^{\prime}Ey^{\prime}\quad\Rightarrow\quad(xEy)\oplus_{B}(x^{\prime\prime}Ey^{\prime\prime})\succsim(x^{\prime}Ey^{\prime})\oplus_{B}(x^{\prime\prime}Ey^{\prime\prime}).$ (I.25)
Theorem 3 (Biseparable Utility Representation [undefad]).